# A Topologically Enriched Probability Monad on the Cartesian Closed Category of CGWH Spaces

### v2

@@ Abstract
Probability monads on categories of topological spaces are classical objects of study in the categorical approach to probability theory. We construct a probability monad on the category of compactly generated weakly Hausdorff (CGWH) spaces, a (if not the) standard choice of convenient category of topological spaces. Because a novel version of the Riesz representation theorem plays a fundamental role in our construction, we name it the Riesz probability monad. We show that the Riesz probability monad is a simultaneous extension of the classical Radon and Giry monads that is topologically enriched. Topological enrichment corresponds to a strengthened continuous mapping theorem (in the sense of probability theory). In addition, restricting the Riesz probability monad to the cartesian closed subcategory of weakly Hausdorff quotients of countably based (QCB) spaces results in a probability monad which is strongly affine, ensuring that the notions of independence and determinism interact as we would expect.
@@ 1.3.2 A Riesz representation theorem
We then prove a Riesz representation theorem (Theorem 4.6),
@@ 1.4 A monads-probability dictionary
Acknowledgements
We gratefully acknowledge Tobias Fritz, for his encouragement, and his helpful comments in the early stages of this work. We would also like to thank Eveliina Peltola for helpful discussions.
2 Compactly generated weakly Hausdorff spaces
In this section, we give an exposition of the category of compactly generated weakly Hausdorff (CGWH) spaces. A thorough account is given in [Str09].
@@ 2 Compactly generated weakly Hausdorff spaces
The following result follows immediately.
@@
The $k$-operator on topologies is not only idempotent but also monotone.
@@ 2.4 The weak Hausdorff quotient
Let $X$ be a CG space. WH quotient of $X$ is $X/\!\sim_{\mathrm{WH}}$, where $\sim_{\mathrm{WH}}$ is the smallest closed (as a subset of $X\times X$) equivalence relation on $X$. (Here, the product is taken in $\mathsf{CG}$, as always.)
@@ 2.6 QCB spaces
sending every point in $X$ to the unique equivalence class of any point $a\in A$. Since equalisers are limits and both $X$ and $X/A$ are weakly Hausdorff QCB spaces, the claim follows from Theorem 2.23. ∎
@@ 2.7 Convergent sequences in mapping spaces
The convergent sequences in a topological space agree with those of its $k$-ification. (This can be seen by identifying convergent sequences with maps from the one-point compactification of the natural numbers.) The topology on $Y^{X}$ is given by the $k$-ification of the compact-open topology This is the topology of uniform convergence on compact subsets, which yields the claim. ∎
@@ 3 Linear CGWH Spaces
CGWH spaces are, by definition, topological spaces and the definition of a linear CGWH space is very similar to that of a topological vector space. However, with respect to the same topology and vector space structure, a linear CGWH space is not necessarily a topological vector space, see [Pet24, Example 2.2.9]. Continuity of addition is only required with respect to the product $\times$ in $\mathsf{CGWH}$, which is in general a weaker requirement than continuity with respect to ×$\times_{\mathsf{Top}}$. In the other direction, a Hausdorff topological vector space is not necessarily a linear CGWH space (with respect to the same topology and vector space structure) either, as exemplified by the weak-* dual of an infinite-dimensional Hilbert space [FR72, Proposition 1].
@@ 3.1 $C_{b}(X)$ as a linear CGWH space
By assumption, $\phi$ is continuous on each space of continuous functions $C_{c.o.}(X,nD)$ with the compact-open topology, implying that it is also continuous on each $(nD)^{X}$ (since the compact-open CG topology is finer than the compact-open topology). The claim follows, since $C_{b}(X)$ is given as the colimit over these spaces. ∎
@@ 3.2 The natural dual of a linear CGWH space
The Cartesian closedness of $\mathsf{CGWH}$ leads to a natural topology on the continuous dual space of a linear CGWH space as follows.
@@
The next lemma will be needed to compare the topology on the space $\mathcal{M}(X)$ of $k$-regular measures (see Section 4.2) to the topology of weak convergence, see Theorem 6.3.
@@
By a corollary of the uniform boundedness principle (see [Sch66, p. 86, 4.6]), the weak-* topology on $V^{\prime}$ coincides with the compact-open topology on the unit ball of $V^{\prime}$. Since the unit ball is compact in the weak-* topology, by the Banach-Alaoglu theorem, it is also compact in the compact-open topology. On compact subsets, the compact-open topology agrees with its $k$-ification, which is precisely the topology of $V^{\wedge}$. ∎
@@ A $k$-regular measure on $X$ is a Baire measure $\mu$ on $X$ for which the integration map,
We consider compact Hausdorff spaces and weakly Hausdorff QCB spaces separately.
@@
We need to show that for every $\phi\in C_{b}(X)^{\wedge}$, there exists a unique Baire measure $\mu$ such that
@@
for all $f\in C_{b}(X)$. By a general version of the Riesz representation theorem [BR07, p. 111, Theorem 7.10.1], this is the case if, and only, if for every monotonically decreasing sequence $(f_{n})$ in $C_{b}(X)$ converging pointwise to zero, $\phi(f_{n})\to 0$ as well. So let $(f_{n})$ be such a sequence. By Dini’s theorem, $(f_{n})$ converges to 00 uniformly on compact subsets of $X$. Moreover, since it is monotonically decreasing, $(f_{n})$ is uniformly bounded by the constant $\sup_{x\in X}f_{0}(x)$. Hence, by Lemma 3.6, $(f_{n})$ converges to 00 in $C_{b}(X)$ and by continuity of $\phi$, $\phi(f_{n})\to 0$, which is what we wanted to show. ∎
As a direct consequence of Lemma 4.5 and Theorem 4.6, we obtain:
@@
We will henceforth identify the set of $k$-regular measures $\mathcal{M}(X)$ with $C_{b}(X)^{\wedge}$. In particular, we topologise $\mathcal{M}(X)$ through this identification. In this way, we obtain an endofunctor,
@@ 5.1 The Riesz monad
Note that under the identification of measures with functionals via Theorem 4.6, we have that
@@
$f_{*}\mu=\mu(-\circ f),\qquad(f:X\to Y,\>\mu\in\mathcal{M}(X),\>X,Y\in\mathsf{CGWH}).$ (2)
@@ 5.2 The Baire monad
5.3 Enriched structure: a strengthened continuous mapping
theorem
A basic fact concerning convergence of random variables in distribution is the continuous mapping theorem, one version of which can be formulated as follows. Let $f:X\to Y$ be a continuous map between metric spaces $X,Y$, and let $(x_{n})$ be a sequence of $X$-valued random variables converging in distribution to an $X$-valued random variable $x$. Then $f(x_{n})$ converges in distribution to $f(x)$. In other words, any weakly convergent sequence of Borel probability measures $(\mu_{n})$ is mapped to a weakly convergent sequence $(f_{*}\mu_{n})$ under the pushforward operation $f_{*}$. When $X,Y$ are Polish spaces we can formulate this even more concisely as the continuity of the pushforward
@@ 5.3 Enriched structure: a strengthened continuous mapping
Let $X,Y$ be complete separable metric spaces, let $(f_{n})$ be a sequence of continuous maps $X\to Y$ converging uniformly on compact sets to $f:X\to Y$, and let $(\mu_{n})$ be a sequence of Borel measures converging weakly to $\mu\in\mathcal{P}(X)$. Then the sequence of pushforwards $((f_{n})_{*}\mu_{n})$ converges weakly to $f_{*}\mu$.
Finally, we note that Proposition 5.4 implies that $\mathcal{M}$ is an enriched monad. Recall that an $\mathsf{V}$-enriched monad on a $\mathsf{V}$-enriched category $\mathsf{C}$ is a $\mathsf{V}$-enriched endofunctor $T$ on $\mathsf{C}$ equipped with two $\mathsf{V}$-enriched natural transformations making the usual diagrams commute. Cartesian closed categories are canonically enriched over themselves, and Proposition 5.4 now implies:
@@ 5.4 Commutativity: product measures and Fubini’s theorem
Let $X,Y$ be QCB spaces. Then the Baire sigma algebra on $X\times Y$ is the product $\sigma$-algebra of the Baire $\sigma$-algebras on each factor,
@@
$\mathcal{B}a(X)\otimes\mathcal{B}a(Y)=\sigma(\{A\times B\,|\,A\in\mathcal{B}a(X),B\in\mathcal{B}a(Y)\}).$
@@
This follows from the fact that QCB spaces are hereditarily Lindelöf (see Proposition 2.24) and [BR07, Proposition 6.10.7]. ∎
@@
$(\mu\otimes\nu)(f)=\nu(y\mapsto\mu(x\mapsto f(x,y))),\qquad(\mu\in\mathcal{M}(X),\nu\in\mathcal{M}(Y),f\in C_{b}(X\times Y)),$
@@
$(\mathcal{M},\delta_{\bullet},\Rbag,\otimes)$ is a symmetric monoidal monad on $\mathsf{QCB}_{h}$.
@@
We need to show that the diagram
@@
commutes (where $\sigma$ is the braiding). Equivalently, we want to show that for all $f\in C_{b}(X)$,
$\int f\>\mathrm{d}\mu\otimes\nu=\int f\>\mathrm{d}\sigma_{*}(\nu\otimes\mu).$
After a change of variables and expanding, this becomes,
@@
which follows from Fubini’s theorem. ∎

### v3

@@ Abstract
Probability monads on categories of topological spaces are classical objects of study in the categorical approach to probability theory, with important applications in the semantics of probabilistic programming languages. We construct a probability monad on the category of compactly generated weakly Hausdorff (CGWH) spaces, a (if not the) standard choice of convenient category of topological spaces. Because a general version of the Riesz representation theorem adapted to this setting plays a fundamental role in our construction, we name it the Riesz probability monad. We show that the Riesz probability monad is a simultaneous extension of the classical Radon and Giry monads that is topologically enriched. Topological enrichment corresponds to a strengthened continuous mapping theorem (in the sense of probability theory). In addition, restricting the Riesz probability monad to the Cartesian closed subcategory of weakly Hausdorff quotients of countably based (QCB) spaces results in a probability monad which is strongly affine, ensuring that the notions of independence and determinism interact as we would expect.
@@ 1.3.2 A Riesz representation theorem
We then prove a Riesz representation theorem (Theorem Representation Theorem),
@@ 1.4 A monads-probability dictionary
2 Compactly generated weakly Hausdorff (CGWH) spaces
In this section, we give an exposition of the category of CGWH spaces. A thorough account is given in [Str09].
@@ 2 Compactly generated weakly Hausdorff (CGWH) spaces
The following result shows that the $k$-operator on topologies is idempotent.
@@
Proof.
If $f:K\rightarrow X$ is $k(\tau)$-continuous, then it is also $\tau$-continuous, because every $\tau$-open set is $k(\tau)$-open.
Now, suppose that $f:K\rightarrow X$ is $\tau$-continuous. It follows immediately from the definition that if $U$ is $k(\tau)$-open, then $f^{-1}(U)$ is open, whence $f$ is $k(\tau)$-continuous. ∎
In addition to idempotency, we also have monotonicity.
@@ 2.1 Limits and colimits of CG spaces
Warning 2.11.
As a particular case of the limits of CG spaces constructed in Lemma 2.9, the product $X\times Y$ of two CG spaces $X$ and $Y$ does not generally coincide with the product $X\times_{\mathsf{Top}}Y$ as taken in the category $\mathsf{Top}$ of topological spaces, see [Eng89, Example 3.3.29]. Similar warnings apply to the various $\sigma$-algebras one may consider on such product spaces; these will play a crucial role later on. A priori, $\mathcal{B}a(X\times Y)$, $\mathcal{B}a(X\times_{\mathsf{Top}}Y)$ and $\mathcal{B}a(X)\otimes\mathcal{B}a(Y)$ might not agree. However, these $\sigma$-algebras do importantly all coincide for weakly Hausdorff QCB spaces (see Lemma 5.7).
@@ 2.4 The weak Hausdorff quotient
Let $X$ be a CG space. The WH quotient of $X$ is $X/\!\sim_{\mathrm{WH}}$, where $\sim_{\mathrm{WH}}$ is the smallest closed (as a subset of $X\times X$) equivalence relation on $X$. (Here, the product is taken in $\mathsf{CG}$, as always.)
@@ 2.6 QCB spaces
sending every point in $X$ to the unique equivalence class of any point $a\in A$. Since equalisers are limits and both $X$ and $X/A$ are weakly Hausdorff QCB spaces, where we use use the fact that $A$ is closed to conclude that $X/A$ is weakly Hausdorff (see [Str09, Corollary 2.21]), the claim follows from Theorem 2.24. ∎
@@ 2.7 Convergent sequences in mapping spaces
The convergent sequences in a topological space agree with those of its $k$-ification. (This can be seen by identifying convergent sequences with maps from the one-point compactification of the natural numbers.) The topology on $Y^{X}$ is given by the $k$-ification of the compact-open topology. This is the topology of uniform convergence on compact subsets, which yields the claim. ∎
@@ 3 Linear CGWH spaces
CGWH spaces are, by definition, topological spaces and the definition of a linear CGWH space is very similar to that of a topological vector space. However, with respect to the same topology and vector space structure, a linear CGWH space is not necessarily a topological vector space [Pet24, Example 2.2.9]. Continuity of addition is only required with respect to the product $\times$ in $\mathsf{CGWH}$, which is in general a weaker requirement than continuity with respect to ×$\times_{\mathsf{Top}}$. In the other direction, a Hausdorff topological vector space is also not necessarily a linear CGWH space with respect to the same topology and vector space structure, as exemplified by the weak-* dual of an infinite dimensional Hilbert space [FR72, Proposition 1].
@@ 3.1 $C_{b}(X)$ as a linear CGWH space
By assumption, $\phi$ is continuous on each space of continuous functions $C_{c.o.}(X,nD)$ with the compact-open topology, implying that it is also continuous on each $(nD)^{X}=kC_{c.o.}(X,nD)$. The claim follows, since $C_{b}(X)$ is given as the colimit over these spaces. ∎
@@ 3.2 The natural dual of a linear CGWH space
The Cartesian closedness of $\mathsf{CGWH}$ suggests a canonical topology on the continuous dual space of a linear CGWH space. When equipped with this topology, we call it the natural dual.
@@
The next lemma will be needed to compare the topology on the space $\mathcal{M}(X)$ of $k$-regular measures (see Section 4.2) to the topology of weak convergence; see Theorem 6.3.
@@
By a corollary of the uniform boundedness principle (see [Sch66, p. 86, 4.6]), the weak-* topology on $V^{\prime}$ coincides with the compact-open topology on the unit ball of $V^{\prime}$. Since the unit ball is compact in the weak-* topology, by the Banach-Alaoglu theorem, it is hence also compact in the compact-open topology. On compact subsets, the compact-open topology agrees with its $k$-ification, which is precisely the topology of $V^{\wedge}$. ∎
@@ A $k$-regular measure on $X$ is a Baire measure $\mu$ on $X$ for which the integration map,
We consider both cases separately.
@@
Assume first that $X$ is Hausdorff. We need to show that for every $\phi\in C_{b}(X)^{\wedge}$, there exists a unique Baire measure $\mu$ such that
@@
for all $f\in C_{b}(X)$. By a general version of the Riesz representation theorem for Hausdorff spaces [BR07, p. 111, Theorem 7.10.1], this is the case if, and only if, for every monotonically decreasing sequence $(f_{n})$ in $C_{b}(X)$ converging pointwise to zero, we have $\phi(f_{n})\to 0$. So let $(f_{n})$ be such a sequence. By Dini’s theorem, $(f_{n})$ converges to 00 uniformly on compact subsets of $X$. Moreover, since it is monotonically decreasing, $(f_{n})$ is uniformly bounded by the constant $\sup_{x\in X}f_{0}(x)$. Hence, by Lemma 3.6, $(f_{n})$ converges to 00 in $C_{b}(X)$ and by continuity of $\phi$, $\phi(f_{n})\to 0$, which completes the proof of the Hausdorff case.
We now turn to the general case, assuming only that $X$ is a weakly Hausdorff CG space. This will follow from applying the Hausdorff case to a certain Hausdorff space $\tilde{X}$ with the property that $C_{b}(X)=C_{b}(\tilde{X})$.
Recall that a topological space $Y$ is functionally Hausdorff if for any two points $y_{1},y_{2}\in Y$ with $f(y_{1})=f(y_{2})$ for all $f\in C(Y)$, we have $y_{1}=y_{2}$. Functionally Hausdorff CG spaces form a reflective subcategory of $\mathsf{CGWH}$. The functionally Hausdorff reflection $\tilde{X}$ of $X$ is given by the quotient of $X$ by the equivalence relation that identifies two points if the values of any real-valued function at both points coincide. Moreover, functionally Hausdorff CG spaces form an exponential ideal in $\mathsf{CGWH}$, i.e. for any CG space $X$ and any functionally Hausdorff space $Y$, we have that $Y^{X}$ is functionally Hausdorff. This implies that $Y^{\tilde{X}}=Y^{X}$, or any CG space $X$ and any functionally Hausdorff space $Y$, and therefore we also have $C(X)=C(\tilde{X})$ and $C_{b}(X)=C_{b}(\tilde{X})$, both as sets and as topological spaces. Hence, a $k$-regular measure on $X$ is equivalently a $k$-regular measure on $\tilde{X}$ and a continuous linear functional on $C_{b}(X)$ is equivalently a continuous linear functional on $C_{b}(\tilde{X})$. Since any functionally Hausdorff space is in particular a Hausdorff space, the general case now follows from the Hausdorff case applied to $\tilde{X}$, as desired. ∎
As a direct consequence of Lemma 4.5 and Theorem Representation Theorem, we obtain:
@@
We will henceforth identify the set of $k$-regular measures $\mathcal{M}(X)$ with $C_{b}(X)^{\wedge}$. In particular, we topologise $\mathcal{M}(X)$ according to the topology on $C_{b}(X)^{\wedge}$. In this way, we obtain an endofunctor,
@@
The next step will be to endow this endofunctor $\mathcal{M}$ with the structure of a monad.
@@ 5.1 The Riesz monad
Note that under the identification of measures with functionals via Theorem Representation Theorem, we have that
@@
$f_{*}\mu=\mu(-\circ f).\qquad(f:X\to Y,\>\mu\in\mathcal{M}(X),\>X,Y\in\mathsf{CGWH})$ (2)
Eq. (1) and (2) display $\mathcal{M}$ as a submonad of the double dualisation monad $\mathbb{K}^{\mathbb{K}^{(-)}}$, which is also know as the continuation monad in the context of programming language theory.
@@ 5.2 The Baire monad
5.3 Enriched structure: a strengthened continuous mapping theorem
A basic fact concerning convergence of random variables in distribution is the continuous mapping theorem [MW43], one version of which can be formulated as follows. Let $f:X\to Y$ be a continuous map between metric spaces $X,Y$, and let $(x_{n})$ be a sequence of $X$-valued random variables converging in distribution to an $X$-valued random variable $x$. Then $f(x_{n})$ converges in distribution to $f(x)$. In other words, any weakly convergent sequence of Borel probability measures $(\mu_{n})$ is mapped to a weakly convergent sequence $(f_{*}\mu_{n})$ under the pushforward operation $f_{*}$. When $X,Y$ are Polish spaces we can formulate this even more concisely as the continuity of the pushforward
@@ 5.3 Enriched structure: a strengthened continuous mapping theorem
Let $X,Y$ be complete separable metric spaces and let $(f_{n})$ be a sequence of continuous maps $X\to Y$ converging uniformly on compact sets to $f:X\to Y$. Moreover, let $(x_{n})$ be a sequence of random variables on $X$ converging in distribution to a random variable $x$ on $X$, i.e. the sequence $(\mu_{n})$ of their distributions converging weakly to the distribution $\mu\in\mathcal{P}(X)$ of $x$. Then $f_{n}(x_{n})$ converges to $f(x)$ in distribution, i.e. the sequence of pushforwards $((f_{n})_{*}\mu_{n})$ converges weakly to $f_{*}\mu$.
Note that, in contrast to the classical continuous mapping theorem, this is a statement about simultaneous convergence in both $(f_{n})$ and $(x_{n})$.
Finally, we remark that Proposition 5.4 implies that $\mathcal{M}$ is an enriched monad. Recall that an $\mathsf{V}$-enriched monad on a $\mathsf{V}$-enriched category $\mathsf{C}$ is a $\mathsf{V}$-enriched endofunctor $T$ on $\mathsf{C}$ equipped with two $\mathsf{V}$-enriched natural transformations making the usual diagrams commute. Cartesian closed categories are canonically enriched over themselves, and Proposition 5.4 now implies:
@@ 5.4 Commutativity: product measures and Fubini’s theorem
Let $X,Y$ be QCB spaces. Then,
@@
$\mathcal{B}a(X)\otimes\mathcal{B}a(Y)=\sigma(\{A\times B\,|\,A\in\mathcal{B}a(X),B\in\mathcal{B}a(Y)\})$
is the product $\sigma$-algebra of the Baire $\sigma$-algebras on each factor.
@@
For the inclusion
$\mathcal{B}a(X)\otimes\mathcal{B}a(Y)\subseteq\mathcal{B}a(X\times Y),$
let $A=f^{-1}(0)$ and $B=g^{-1}(0)$ with $f\in C(X),g\in C(Y)$. Since $\mathcal{B}a(X)\otimes\mathcal{B}a(Y)$ is generated by cartesian products of the form $A\times B$, it suffices that $A\times B\in\mathcal{B}a(X\times Y)$, which holds, as $A\times B=(f\cdot g)^{-1}(0)$.
For the reverse inclusion, it suffices to show that every continuous map $X\times Y\to\mathbb{R}$ is measurable with respect to the product $\sigma$-algebra $\mathcal{B}a(X)\otimes\mathcal{B}a(Y)$. We will show the stronger claim that, in fact, every separately continuous function $X\times Y\to\mathbb{R}$ is $\mathcal{B}a(X)\otimes\mathcal{B}a(Y)$-measurable. In order to achieve this, we will apply certain point-set topological results regarding so-called Rudin spaces.
A topological space $Z$ is said to be a Rudin space if every separately continuous function on $Z\times W$, where $W$ is any other topological space, is the pointwise limit of a sequence of continuous functions on $Z\times_{\mathsf{Top}}W$. We will show that every QCB space is a Rudin space. This implies the original claim, since every continuous function on $X\times_{\mathsf{Top}}Y$ is $\mathcal{B}a(X)\otimes\mathcal{B}a(Y)$-measurable (by [BR07, Proposition 6.10.7] together with the fact that QCB spaces are hereditarily Lindelöf, see Proposition 2.25) and pointwise limits of measurable functions are measurable.
Let $Z$ be any QCB space. By [Ban08, Theorem 6.2.(2)], $Z$ is a Rudin space if $C_{p}(Z)$ is a Rudin space, where $C_{p}(Z)$ is the space of continuous functions on $Z$ with the pointwise topology (i.e. the topology that it inherits from the $\mathsf{Top}$-product of $|Z|$-fold many copies of the real line). Now, $C_{p}(Z)$ has a countable network, since it is the continuous image of the QCB space $C(Z)$ under the identity map. Moreover, as a topological vector space, $C_{p}(Z)$ is normal. Hence, by the Lindelöf property, $C_{p}(Z)$ is paracompact. As a paracompact space with a countable (in particular, $\sigma$-discrete) network, by [Ban08, Theorem 6.2.(3), 2.3.(1)], $C_{p}(Z)$ is Rudin space and hence, so is $Z$. This shows that every QCB space is a Rudin space, which completes the proof. ∎
@@
$(\mu\otimes\nu)(f)=\nu(y\mapsto\mu(x\mapsto f(x,y))),\quad(\mu\in\mathcal{M}(X),\nu\in\mathcal{M}(Y),f\in C_{b}(X\times Y)),$
@@
$(\mathcal{M},\delta_{\bullet},\Rbag,\otimes)$ is a symmetric monoidal monad on $\mathsf{QCB}_{h}$. Equivalently [Koc72], $\mathcal{M}$ is a commutative strong monad with respect to the associated left strength,
$\lambda^{X,Y}:X\times\mathcal{M}(Y)\to\mathcal{M}(X\times Y),\;\;(x,\nu)\mapsto\delta_{x}\otimes\nu,$
and right strength,
$\rho^{X,Y}:\mathcal{M}(X)\times Y\to\mathcal{M}(X\times Y),\;\;(\mu,y)\mapsto\mu\otimes\delta_{y}.$
@@
We need to show that the following diagram commutes:
@@
Let $\mu_{0}\in\mathcal{M}(x),\nu_{0}\in\mathcal{M}(Y)$ and $f\in C_{b}(X\times Y)$. Then,
$(\Rbag\circ\mathcal{M}(\rho)\circ\lambda(\mu_{0},\nu_{0}))(f)$ $=\;[\Rbag(\mathcal{M}(\rho)(\delta_{\mu_{0}}\otimes\nu_{0}))](f)$
$=\;[\Rbag(F\mapsto(\delta_{\mu_{0}}\otimes\nu_{0})[(\mu,y)\mapsto F(\mu\otimes\delta_{y})])](f)$ (Eq. (2))
$=\;[\Rbag(F\mapsto\nu_{0}(y\mapsto F(\mu_{0}\otimes\delta_{y})])](f)$
$=\nu_{0}(y\mapsto(\mu_{0}\otimes\delta_{y})(f))$ (Eq. (1))
$=\nu_{0}(y\mapsto\mu_{0}(x\mapsto f(x,y)))$
$=\int_{Y}\int_{X}f(x,y)\mathrm{d}\mu_{0}(x)\mathrm{d}\nu_{0}(y).$
By symmetry, we also have,
$(\Rbag\circ\mathcal{M}(\lambda)\circ\rho(\mu_{0},\nu_{0}))(f)=\int_{X}\int_{Y}f(x,y)\mathrm{d}\nu_{0}(y)\mathrm{d}\mu_{0}(x).$
Hence, our claim reduces to,
@@
which follows from Fubini’s theorem, using Lemma 5.7. ∎